# Large Multimodal Models for Embodied Intelligent Driving: The Next Frontier in Self-Driving?

### v2

@@ Abstract
The advent of Large Multimodal Models (LMMs) offers a promising technology to tackle the limitations of modular design in autonomous driving, which often falters in open-world scenarios requiring sustained environmental understanding and logical reasoning. Besides, embodied artificial intelligence facilitates policy optimization through closed-loop interactions to achieve the continuous learning capability, thereby advancing autonomous driving toward embodied intelligent (El) driving. However, such capability will be constrained by relying solely on LMMs to enhance EI driving without joint decision-making. This article introduces a novel semantics and policy dual-driven hybrid decision framework to tackle this challenge, ensuring continuous learning and joint decision. The framework merges LMMs for semantic understanding and cognitive representation, and deep reinforcement learning (DRL) for real-time policy optimization. We starts by introducing the foundational principles of EI driving and LMMs. Moreover, we examine the emerging opportunities this framework enables, encompassing potential benefits and representative use cases. A case study is conducted experimentally to validate the performance superiority of our framework in completing lane-change planning task. Finally, several future research directions to empower EI driving are identified to guide subsequent work.
@@ I Introduction
In this context, embodied AI has emerged as a revolutionary means of promoting autonomy in autonomous driving systems. Unlike conventional AI models, embodied AI emphasizes the interactions between embodied agents and their physical environment [8]. By embedding embodied AI, self-driving vehicle can adapt to dynamic environmental changes and make intelligent decisions based on the interactive data. Such adaptability enables the vehicles to continuously optimize and evolve their driving policies through interaction in complex environment, thereby realizing the embodied intelligent (EI) driving paradigm. Currently, embodied AI is in the exploratory stage, used for real-time decision-making and action planning in dynamic and uncertain environment. Great potential has been particularly demonstrated in the transportation-related scenarios, including vehicular networks and rail transportation systems [9, 10, 11, 12].
@@
Although these achievements are of great significance, as shown in Table I, a closed-loop design mechanism capable of continuous learning in response to environmental changes within dynamic traffic conditions has yet to be established. While LMMs enhance understanding of multimodal driving data, deep reinforcement learning (DRL) is required to facilitate joint decision-making in dynamic environment. The decision process often necessitates continuous interaction between EI vehicle and its environment. Therefore, incorporating the continuous learning and joint decision in dynamic scenarios into EI driving remains an emergent research frontier.
In this article, we propose employing the joint decision based on LMMs and DRL to enhance EI driving. LMMs inherit and extend the capabilities of LLMs, traditionally designed for textual data processing. Through architectural adaptation and multimodal collaborative training via LMMs, EI vehicle effectively handles the collected multimodal data [13]. These data processed and integrated into internal representation thereby enhance its generalization ability to new environment and tasks. In contrast, DRL behaviors can be adjusted by real-time feedback, which results in adaptive policy development in dynamic operations [14, 15]. In DRL, the agent interacts with the environment and as a result, receives feedback in the form of rewards or penalties. And thus, the agent learns how to make optimal decisions. We will now demonstrate the application of LMMs and DRL-based joint decision across various scenarios, as shown in Fig. 1. The contributions of this article are summarized below.
@@
Figure 1: An illustration of the EI driving scenarios enabled by the joint decision approach based on LMMs and DRL. Such approach shows great promise in three illustrative application scenarios: i) Urban robotaxi, which is on-demand and driverless transport, faces tremendous challenges in negotiating highly dynamic traffic scenarios. Examples would be predicting a pedestrians intent, and avoiding a sudden obstacle in a narrow corridor. ii) Autonomous freight operates in trunk-line logistics, ports, and industrial parks, where good performance is needed in complex geometrical road situations. Representative scenarios include container handling, and narrow space delivery robots. iii) Structured shuttle, driving in structured and semi-open environments, must achieve precision docking. Typical scenarios include airport shuttle service, and micro-circulation bus.
@@ II An Overview of LMMs and EI Driving
This section begins with the fundamental concept of EI driving, followed by the general framework of LMMs. Then, the necessities of integrating EI driving and LMMs is discussed.
@@ II-B Fundamentals of LMMs
Multimodal Instruction Tuning. Multimodal instruction tuning aims to comprehend and carry out cross-modal instructions, which refers to building the pairs of instruction semantics to the output data. LMMs applies the In-Context Learning (ICL) and Chain-of-Thought (CoT) reasoning capability in a multimodal environment, enhancing their cross-modal generalization performance. In complex and varied cross-modal tasks, performance enhancement is especially observed.
@@ III-A Framework
Stage 2. Policy Pipeline. To ensure real-time performance optimization of EI vehicle in dynamic and open environment, the policy pipeline employs DRL approach that directly maps the collected multimodal data into executable action policy. Within this DRL framework, the policy network parameters are continuously optimized through online interaction mechanism [1]. These updates strategically integrate the recommend actions generated by semantic reasoning as the policy references for joint decision, thereby generating the execution action policy. This ensures the dynamic responsiveness to evolving environmental states, achieving stable and operationally efficient control behaviors. The DRL framework is described in detail below.
@@ IV A Case Study of Lane-Change Planning
Dataset. Throughout the experiments, we employ the open-source nuScenes dataset [^2] https://www.nuscenes.org/nuscenes., which integrates the structured data from three modalities. These modalities include multi-view scene videos (local), bird’s-eye-view (BEV) map images (global), and multi-round Question Answering (QA) annotations. The nuScenes dataset consists of images from 1,000 driving scenes, captured across diverse urban and highway environments. Besides, it includes 18,000 reasoning-based QA pairs, and more than 50,000 reasoning steps. This ensures the model learns to generate structured rationales before arriving at a final answer. We utilize a subset of this dataset as the multimodal information input for LMMs. These data are used for the preliminary training of EI vehicle’s lane-change task.
@@ V-B Endogenous Security Defense
With the popularization of V2X communications, EI driving system is increasingly exposed to multiple cyber threats. Existing security mechanisms often struggle to promptly respond to such emerging risks. Endogenous security is the embedding of adaptive and self-evolving security solutions in the core architecture of the system with a security consideration from the design stage itself. This approach establishes the proactive defense paradigm, enabling the system to autonomously identify, resist, and mitigate various security threats. Therefore, integrating such security‑endogenous paradigm into the system architecture represent a promising research direction.

### v3

@@ Abstract
The advent of Large Multimodal Models (LMMs) offers a promising technology to tackle the limitations of modular design in autonomous driving, which often falters in open-world scenarios requiring sustained environmental understanding and logical reasoning. Besides, embodied artificial intelligence facilitates policy optimization through closed-loop interactions to achieve the continuous learning capability, thereby advancing autonomous driving toward embodied intelligent (El) driving. However, such capability will be constrained by relying solely on LMMs to enhance EI driving without joint decision-making. This article introduces a novel semantics and policy dual-driven hybrid decision framework to tackle this challenge, ensuring continuous learning and joint decision. The framework merges LMMs for semantic understanding and cognitive representation, and deep reinforcement learning (DRL) for real-time policy optimization. We start by introducing the foundational principles of EI driving and LMMs. Moreover, we examine the emerging opportunities this framework enables, encompassing potential benefits and representative use cases. A case study is conducted experimentally to validate the performance superiority of our framework in completing lane-change planning task. Finally, several future research directions to empower EI driving are identified to guide subsequent work.
@@ I Introduction
In this context, embodied AI has emerged as a revolutionary means of promoting autonomy in autonomous driving systems. Unlike conventional AI models, embodied AI emphasizes the interactions between embodied agents and their physical environment [8]. By embedding embodied AI, self-driving vehicles can adapt to dynamic environmental changes and make intelligent decisions based on the interactive data. Such adaptability enables the vehicles to continuously optimize and evolve their driving policies through interaction in complex environments, thereby realizing the embodied intelligent (EI) driving paradigm. Currently, embodied AI is in the exploratory stage, used for real-time decision-making and action planning in dynamic and uncertain environments. Great potential has been particularly demonstrated in the transportation-related scenarios, including vehicular networks and rail transportation systems [9, 10, 11, 12].
@@
Although these achievements are of great significance, as shown in Table I, a closed-loop design mechanism capable of continuous learning in response to environmental changes within dynamic traffic conditions has yet to be established. While LMMs enhance understanding of multimodal driving data, deep reinforcement learning (DRL) is required to facilitate joint decision-making in dynamic environments. The decision process often necessitates continuous interaction between EI vehicles and their environment. Therefore, incorporating the continuous learning and joint decision in dynamic scenarios into EI driving remains an emergent research frontier.
In this article, we propose employing the joint decision based on LMMs and DRL to enhance EI driving. LMMs inherit and extend the capabilities of LLMs, traditionally designed for textual data processing. Through architectural adaptation and multimodal collaborative training via LMMs, EI vehicles effectively handle the collected multimodal data [13]. These data processed and integrated into internal representation thereby enhance its generalization ability to new environments and tasks. In contrast, DRL behaviors can be adjusted by real-time feedback, which results in adaptive policy development in dynamic operations [14, 15]. In DRL, the agent interacts with the environment and as a result, receives feedback in the form of rewards or penalties. And thus, the agent learns how to make optimal decisions. We will now demonstrate the application of LMMs and DRL-based joint decision across various scenarios, as shown in Fig. 1. The contributions of this article are summarized below.
@@
Figure 1: An illustration of the EI driving scenarios enabled by the joint decision approach based on LMMs and DRL. Such approach shows great promise in three illustrative application scenarios: i) Urban robotaxi, which is on-demand and driverless transport, faces tremendous challenges in negotiating highly dynamic traffic scenarios. Examples would be predicting a pedestrian’s intent, and avoiding a sudden obstacle in a narrow corridor. ii) Autonomous freight operates in trunk-line logistics, ports, and industrial parks, where good performance is needed in complex geometrical road situations. Representative scenarios include container handling, and narrow space delivery robots. iii) Structured shuttle, driving in structured and semi-open environments, must achieve precision docking. Typical scenarios include airport shuttle service, and micro-circulation bus.
@@ II An Overview of LMMs and EI Driving
This section begins with the fundamental concept of EI driving, followed by the general framework of LMMs. Then, the necessities of integrating EI driving and LMMs are discussed.
@@ II-B Fundamentals of LMMs
Multimodal Instruction Tuning. Multimodal instruction tuning aims to comprehend and carry out cross-modal instructions, which refers to building the pairs of instruction semantics to the output data. LMMs apply the In-Context Learning (ICL) and Chain-of-Thought (CoT) reasoning capability in a multimodal environment, enhancing their cross-modal generalization performance. In complex and varied cross-modal tasks, performance enhancement is especially observed.
@@ III-A Framework
Stage 2. Policy Pipeline. To ensure real-time performance optimization of EI vehicle in dynamic and open environment, the policy pipeline employs DRL approach that directly maps the collected multimodal data into executable action policy. Within this DRL framework, the policy network parameters are continuously optimized through online interaction mechanism [1]. These updates strategically integrate the recommended actions generated by semantic reasoning as the policy references for joint decision, thereby generating the execution action policy. This ensures the dynamic responsiveness to evolving environmental states, achieving stable and operationally efficient control behaviors. The DRL framework is described in detail below.
@@ IV A Case Study of Lane-Change Planning
Dataset. Throughout the experiments, we employ the open-source nuScenes dataset [^2] https://www.nuscenes.org/nuscenes., which integrates the structured data from three modalities. These modalities include multi-view scene videos (local), bird’s-eye-view (BEV) map images (global), and multi-round Question Answering (QA) annotations. The nuScenes dataset consists of images from 1,000 driving scenes, captured across diverse urban and highway environments. Besides, it includes 18,000 reasoning-based Question Answering pairs, and more than 50,000 reasoning steps. This ensures the model learns to generate structured rationales before arriving at a final answer. We utilize a subset of this dataset as the multimodal information input for LMMs. These data are used for the preliminary training of EI vehicle’s lane-change task.
@@ V-B Endogenous Security Defense
With the popularization of V2X communications, EI driving system is increasingly exposed to multiple cyber threats. Existing security mechanisms often struggle to promptly respond to such emerging risks. Endogenous security is the embedding of adaptive and self-evolving security solutions in the core architecture of the system with a security consideration from the design stage itself. This approach establishes the proactive defense paradigm, enabling the system to autonomously identify, resist, and mitigate various security threats. Therefore, integrating such security‑endogenous paradigm into the system architecture represents a promising research direction.